• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 16:02:44 on April21,2008

    Entry number 233611

    keyword=Day shift summary

    Day shift summary.
    Shift workers: J. P. Chen (SL), R. Lindgren (TO), M. Mihovilovic (3rd).

    8:00 production data taking with 4.7 uA at kin. H.
    8:25 Hall C stopped production, starting Moller.
    8:50 ROC5 crash and Chamber HVs (L14.5, L14.6) tripped, reset.
    (end run 3564). Run 3565 started when chamber HVs still being reset.
    9:40 beam shot up whne recovering from trip.
    9:55, beam trip fruqently.
    10:40 coda stuck (run 3569), reboot and restarted coda.
    11:09, tunred on raster: asked for 0.5 x 0.7, MCC could not get that small. After a few tries, they got us to 0.85x0.85mm, give us (spot++) 0.6x1mm. runs 3571-3573.
    11:25 beam current increased to 6.6 uA (run 3574).
    12:10 beam off
    12:20 bem back, beam current to 5.5 uA (run 3577).
    14:02 requested beam current to 4 uA (run 3582).
    14:50 MCC called that they have RF problem, beam off.
    15:28 beam back, requested 5 uA.
    15:55 completed data taking for LH2 at kin H (0.356C).
    move target to 6cm dummy
    15:59 5 uA on dummy run 3585.

    Kin. H., 4.7 uA on 6 cm LH2.
    Run 3562/3563 good runs,
    3564, ROC5 crash/chamber HVs trip.
    3565 started when chamber HVs still being reset.
    3566/3567/3568 good runs.
    3569 coda stuck.
    3570 good run.
    3571-3573 raster checks.
    3574/3575/3576 production with 6.6 uA with raster on.
    3577/3578/3579/3580/3581 beam current 5.4 uA.
    3582/3583 beam current 3.9 uA.
    3584 beam current 4.9 uA.

    5 uA on 6 cm dummy at kin H
    3585. still on-gonig at end of shift.